### v2

# Fact-Checking at Scale: Multimodal AI for Authenticity and Context Verification in Online Media

Abstract.
The proliferation of multimedia content on social media platforms has dramatically transformed how information is consumed and disseminated. While this shift enables real-time coverage of global events, it also facilitates the rapid spread of misinformation and disinformation, especially during crises such as wars, natural disasters, or elections. The rise of synthetic media and the reuse of authentic content in misleading contexts have intensified the need for robust multimedia verification tools. In this paper, we present a comprehensive system developed for the ACM Multimedia 2025 Grand Challenge on Multimedia Verification. Our system assesses the authenticity and contextual accuracy of multimedia content in multilingual settings and generates both expert-oriented verification reports and accessible summaries for the general public. We introduce a unified verification pipeline that integrates visual forensics, textual analysis, and multimodal reasoning, and propose a hybrid approach to detect out-of-context (OOC) media through semantic similarity, temporal alignment, and geolocation cues. Extensive evaluations on the Grand Challenge benchmark demonstrate the system’s effectiveness across diverse real-world scenarios. Our contributions advance the state of the art in multimedia verification and offer practical tools for journalists, fact-checkers, and researchers confronting information integrity challenges in the digital age.
Multimedia Verification, Fact-Checking, Multi-Agent Systems, 2qSzLarge Language Models (LLM)
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Computer visionccs: Information systems Multimedia information systems
1. Introduction
In the digital age, social media platforms have become a primary channel for information consumption and dissemination. An enormous volume of multimedia content, including images, videos, and user-generated posts, is uploaded every day across platforms such as Twitter, Facebook, TikTok, and Telegram. This rapid circulation of visual content enables real-time coverage of global events but also creates fertile ground for the spread of misinformation and disinformation. Studies have shown that false content spreads significantly faster and reaches more users than accurate information, particularly during high-impact events like wars, disasters, or elections (Vosoughi et al., 2018; Gupta et al., 2013). The use of synthetic media, including deepfakes and out-of-context visual narratives, has further amplified concerns about public trust and information reliability.
To address these risks, multimedia verification has become a critical aspect of modern fact-checking workflows. The task involves assessing the authenticity, origin, context, and integrity of multimedia content, often using multiple sources and modalities. Recent progress in Open Source Intelligence (OSINT) has supported this mission through a growing ecosystem of tools and platforms designed for investigative journalism and digital forensics. Prominent examples include Bellingcat’s Online Investigation Toolkit (bel, [n. d.]), Google Fact Check Explorer (goo, [n. d.]), and the OSINT Framework (osi, [n. d.]). Despite these efforts, several key challenges persist. Multimedia verification requires handling diverse modalities (e.g., image, video, text, and metadata), detecting subtle manipulations, and reasoning about context and source credibility.
In this paper, we introduce a complete system developed for the ACM Multimedia 2025 Grand Challenge on Multimedia Verification (Dang-Nguyen et al., 2025). Our system is designed to evaluate both the authenticity and contextual accuracy of multimedia content in multilingual settings. It outputs a structured, human-readable verification report for fact-checkers, as well as a simplified summary for broader consumption. Our solution also addresses the Out-of-Context (OOC) sub-task, identifying misleading or de-contextualized uses of otherwise authentic media.
Our main contributions are as follows:
  • We develop a unified multimedia verification pipeline that integrates visual forensics, text analysis, and multilingual support.
  • We propose an interpretable reporting structure tailored to expert fact-checkers and general users alike.
  • We introduce a hybrid approach for OOC detection that combines semantic similarity, temporal alignment, and geolocation cues.
  • We validate the effectiveness of our system through participation in the ACM Multimedia 2025 Grand Challenge on Multimedia Verification.
The rest of the paper is structured to answer the key questions posed by the Grand Challenge and to demonstrate the practical utility of our system across real-world verification scenarios.
2. Proposed Method
The multimedia verification framework employs a multi-stage pipeline designed to comprehensively analyze multimedia content through evidence aggregation and contextual analysis. The system integrates four core verification services that address the fundamental questions of multimedia authenticity: geolocation (”Where?”), temporal analysis (”When?”), content classification, and intent analysis (”Why?”). The framework processes diverse input modalities including JSON metadata, images, and videos, converting them into standardized formats for downstream analysis services.
The overall architecture follows a unified approach that combines visual forensics, natural language processing, and web-based evidence gathering to produce structured verification reports. The system leverages large language models for content classification and reasoning, while employing specialized deep learning models for geolocation and temporal analysis. Each verification service operates independently but contributes to a comprehensive evidence aggregation module that synthesizes findings into interpretable reports for fact-checkers and end users.
Figure 1. The pipeline begins with three types of inputs: JSON metadata, images, and videos. In the data preprocessing stage, metadata is processed via a social media crawler, transcripts are extracted from video content using OpenAI Whisper, while images and video frames—extracted using Google Vision Intelligent keyframe extraction—are encoded in Base64 format for standardized handling. The preprocessed data is then routed to four specialized verification services: ”Where?” (geolocation), ”When?” (temporal analysis), content classification, and AIGVDet (AI-generated visual detection). Outputs from these services are aggregated in the evidence aggregation module, which performs source verification, intent analysis (”Why?”), and report generation. The final output is a structured verification report designed to support fact-checking and multimedia forensics.
2.1. Data Preprocessing
The data preprocessing pipeline handles heterogeneous multimedia inputs through a systematic three-stage process: metadata extraction, media processing, and format standardization.
  • Metadata Processing: The system processes JSON metadata files containing structured information including title, description, location, category, violence level, and social media links. This metadata serves as contextual information for subsequent verification services and provides initial cues for content classification and temporal analysis.
  • Media Encoding: All processed media content undergoes Base64 encoding to enable standardized API transmission across verification services. The system implements separate encoding functions for static images and video frames, ensuring consistent format handling regardless of input type.
  • Integration and Standardization: The preprocessing module integrates metadata and processed media into a unified data structure compatible with downstream verification services. This includes organizing media content, extracting relevant textual features, and preparing data payloads for geolocation analysis, temporal verification, and content classification services. The standardized output format enables seamless integration with the evidence aggregation pipeline while maintaining data integrity throughout the verification.
2.2. Content Classification
Content classification is a crucial step in organizing and understanding the verified multimedia content. This process aims to categorize the content by assigning relevant tags based on its subject matter, origin, and key entities involved. Large Language Models (LLMs) are employed for this task, leveraging their advanced semantic understanding to generate a comprehensive set of keywords. These tags are derived from the aggregated textual evidence gathered from supporting sources. The generated tags span various categories, including the social media platforms where the content circulated (e.g., TikTok, Telegram, X), prominent individuals or organizations depicted or mentioned (e.g., Trump, Coca-Cola, Hamas, Al Jazeera), and specific events or topics to which the content pertains (e.g., Ukraine War, AI-generated, Gaza, Shifa Hospital). This systematic tagging facilitates efficient retrieval and contextualization of verified information.
2.3. Verified Evidence
2.3.1. Source Details
By analyzing the gathered evidence using LLM, the system meticulously identifies key attributes for each supporting source. This includes pinpointing the original publisher or entity responsible for the content, specifying the platform on which it appeared (e.g., social media, news website), and providing direct links (URLs) to the original posts or articles. Additionally, any available publication dates and timestamps are extracted. The aim is to present a clear, traceable lineage of the content, enabling fact-checkers to assess its primary distribution and initial context effectively.
2.3.2. Where? (Location)
To estimate the geographic locations of images and videos, an enhanced version of the G3 framework (Jia et al., 2024) is adopted, as illustrated in Figure 2. The pipeline comprises two main phases: Database Construction and Location Prediction.
Phase 1: Database Construction. This phase builds an image database using the pretrained model obtained from the Geo-alignment process, in order to support the large multimodal model (LMM) in the next phase.
  • Geo-alignment. In this step, images, textual descriptions, and GPS coordinates are aligned within a shared embedding space. Visual and textual features are extracted using frozen CLIP encoders, while GPS coordinates are processed through a Mercator projection followed by Random Fourier Feature encoding. All representations are jointly trained using contrastive loss.
  • Image Database. The database is constructed from the MP16-Pro dataset (Jia et al., 2024). Each image is embedded using the pretrained image encoder, and the resulting visual features are combined with corresponding textual and GPS-aligned representations to retain geolocation information.
Phase 2: Location Prediction. To estimate the final GPS coordinates and location, two key steps from G3 (Jia et al., 2024) are used: Geo-diversification and Geo-verification.
  • Geo-diversification. The system leverages Google’s Gemini Pro 2.5 as the LMM for a Retrieval-Augmented Generation (RAG) pipeline. To enable LMM to reason over rich contextual information, the input prompt includes similar and dissimilar GPS coordinates, metadata, available transcripts, and search results retrieved from Cloud Vision Web Detection and Scrapingdog. The LMM processes multiple such prompts to generate candidate predictions, each consisting of a GPS estimate, the corresponding location, and supporting evidence.
  • Geo-verification. This process starts by encoding each predicted GPS and computing their cosine similarity against the input images. The prediction with the highest average similarity score is then selected, and its location is verified by the LMM and OpenStreetMap’s Nominatim service. Finally, the supporting evidence is augmented with images of the predicted location, ensuring that each piece of evidence is supported by input images, relevant location images, and URLs.
Figure 2. The pipeline of finding (A) where and (B) when images/ videos are taken
2.3.3. When? (Time)
To detect the capture or recording time of multimedia content (images/videos), we propose a practical pipeline that leverages metadata, vision-language modeling, and web search engines.
Given an input multimedia file $M$ (image or video) and an associated metadata file $J$ (in JSON format), the system outputs:
  • Timestamp: The estimated capture or recording time of $M$.
  • Source: A relevant webpage or article containing $M$ or similar visual evidence.
  • Confidence: The similarity score between metadata and textual content of the source.
Two types of search are performed:
  • Textual Search: A search query is constructed from the metadata:
    $q=\texttt{location}+\texttt{title}+\texttt{description}$
    and submitted to Google Web Search API via SerpAPI.
  • Reverse Image Search: For each keyframe (or the image input), the image is uploaded to imgbb to obtain a public URL, which is then used in Google Reverse Image Search via SerpAPI.
The results of both search types are merged.
Each result is scored using:
  • Textual similarity: Using SequenceMatcher between the result’s title+snippet and the metadata query $q$.
  • Timestamp parsing: Dates are extracted and normalized to standard formats.
Results are sorted in descending order of similarity. Ties are broken using ascending publication date. Due to cost constraints, the system adopts an early stopping strategy: the first high-confidence result containing a valid timestamp terminates the search process.
To comply with the limited query quota of SerpAPI (100 searches /month on the free tier), we employ an early stopping mechanism that halts the search once a confident timestamp is found. Although this reduces cost, it may overlook stronger evidence in later keyframes. A fully exhaustive approach—processing all keyframes and ranking all results—would improve reliability but requires more query budget. Future improvements may include adaptive stopping or cost-aware scoring policies.
2.3.4. Why? (Motivation or Intent)
In our methodology, we leverage the enhanced reasoning and explainability capabilities of large language models (LLMs), facilitated by advanced prompting strategies such as chain-of-thought. Specifically, our system constructs a reasoned explanation of the model’s intent by aggregating intermediate evidence generated from earlier processing stages and applying multi-level reasoning. This structured reasoning process supports the derivation and justification of the final decision output.
Evidences Gathering
The gathering of evidence systematically acquires relevant external information to augment the multimedia verification process. The approach begins with a reverse image search that utilizes the input media (or extracted keyframes from video) to identify corresponding and partially matching web pages. The URLs thus identified are subsequently processed to extract contextual information. For standard Web articles, core content, titles, authors, and publication dates are extracted via an automated parsing mechanism. Specialized handling is used for social media platforms, such as X (formerly Twitter), where post content and metadata are retrieved through a dedicated scraping approach. All external links collected and their corresponding textual content are aggregated, establishing a comprehensive pool of supporting evidence for subsequent analysis stages, including content classification and intent analysis.
2.4. Forensic Analysis
To assess video authenticity, we adopt the AIGVDet model (Bai et al., 2024), a forensic framework designed to detect AI-generated content. AIGVDet comprises two key modules: a Spatial Domain Detector and an Optical Flow Detector.
The Spatial Domain Detector analyzes individual RGB frames to identify visual artifacts such as unnatural textures, lighting inconsistencies, and anomalous pixel patterns. Features are extracted using a ResNet50 backbone (He et al., 2016). In parallel, the Optical Flow Detector captures temporal artifacts using flow maps computed via the RAFT algorithm (Teed and Deng, 2020), exploiting the fact that generative models often fail to preserve coherent motion.
Let $I_{i}$ denote the $i$-th frame, and $F_{i}=\mathcal{F}(I_{i},I_{i+1})$ the corresponding optical flow. Features $V_{i}^{I}$ and $V_{i}^{F}$ are extracted using ResNet50 encoders $R(\cdot)$ and $O(\cdot)$:
(1) $V_{i}^{I}=R(I_{i}),\quad V_{i}^{F}=O(F_{i})$
These features are passed through a global average pooling layer and a fully connected layer with sigmoid activation:
(2) $P_{i}^{I}=\sigma(\mathrm{FC}(\mathrm{GAP}(V_{i}^{I}))),\quad P_{i}^{F}=\sigma(\mathrm{FC}(\mathrm{GAP}(V_{i}^{F})))$
The fused frame-level prediction is:
(3) $P_{i}=\alpha P_{i}^{I}+(1-\alpha)P_{i}^{F}$
The final video-level score is the average across all frames:
(4) $P=\frac{1}{N-1}\sum_{i=1}^{N-1}P_{i}$
AIGVDet is trained on the Generated Video Dataset (GVD), which includes synthetic videos from 11 generative models spanning both text-to-video and image-to-video pipelines (Bai et al., 2024). In our pipeline, we adopt AIGVDet not for binary classification, but to produce confidence scores for both authentic and synthetic labels. This probabilistic output provides a more nuanced and interpretable assessment of video authenticity.
2.5. Summary of Key Points
Our system leverages Large Language Models (LLMs) to synthesize a concise and comprehensive case summary from all available verification evidence. This includes initial textual supporting sources, contextual media information, and detailed findings from the ”When?”, ”Where?”, ”Who?”, and ”Why?” analysis modules. The LLM is guided by a structured prompt and a defined JSON schema to ensure the output is consistent, accurate, and readily interpretable. The generated summary integrates findings across all modalities, highlighting the central claim or event, primary verification outcomes, and any identified uncertainties or unknowns.
2.6. Sub-Task: OOC Detection
2.6.1. SearchOOC
We propose a method called SearchOOC for detecting whether an image-caption pair is out-of-context (OOC) or not (Figure 3):
  1. (1)
    Search for Evidence: We use Google Search with the input image-caption pair to obtain a list of URLs that likely contain both the image and the caption.
  2. (2)
    Check for OOC: Among the retrieved URLs, we check whether any site explicitly mentions that the pair is out-of-context. If we find a website making this determination, SearchOOC classifies the input pair as OOC. Otherwise, we proceed to the next step.
  3. (3)
    Check for NOOC: We check whether any of the URLs belong to a well-known website that restricts access (such as The New York Times). If so, we classify the input pair as not out-of-context (NOOC). If not, we further verify if any website actually contains the input image and caption. Since the image might be resized and the caption paraphrased, we use image hashing (with a threshold of 10) and text similarity (with a threshold of 7.5) to make this determination. If either check passes, we classify the pair as NOOC. If not, we move to the next step.
  4. (4)
    Detect Using an Alternative Method: In this step, we apply a baseline model (o4-mini) with the following direct prompt: “Given an image and the caption, determine whether the image and caption are out of context. Please output only ‘ooc’ or ‘nooc’ without any explanation.”
Figure 3. SearchOOC for detecting out-of-context
2.6.2. HierOOC
HierOOC is a hierarchical, multi-stage framework designed to detect whether a given input—including an image, a caption, and optionally a context—is out-of-context (OOC). The system consists of four iterative stages:
  1. (1)
    Multi-scale Evidence Retrieval: Given the input pair, this stage retrieves evidence at two levels. At the external scale, a search engine is used to collect relevant URLs and news articles. At the internal scale, an entity detector extracts textual and visual entities from the image-caption pair.
  2. (2)
    Evidence Filtering: To ensure evidence quality and relevance, HierOOC applies several filtering strategies, including domain filtering, language filtering, redundancy removal, and similarity-based filtering. A similarity threshold of 0.7 is used to retain only the most relevant candidates.
  3. (3)
    Evidence Refinement: Some retrieved evidence, particularly from web sources, may still be inconsistent with the original input and internal candidates. Therefore, a large language model (LLM) is employed to assess and eliminate inconsistencies, thereby reducing the input token size for subsequent stages instead of passing all collected evidence forward.
  4. (4)
    OOC Decision Making: In the final stage, the OOC Decisor determines whether the input is out-of-context. It takes the input pair and the refined evidence summary as input. This module includes detection agents guided by Chain-of-Thought (CoT) prompting to improve reasoning and reduce hallucination. All LLMs used in this framework are based on the Gemini 2.5 Flash Lite version.
Figure 4. HierOOC for detecting out-of-context
3. Experimental Result
3.1. Dataset
3.1.1. Main-task
The main-task dataset contains real-world multimedia cases, each with images or videos, multilingual context (such as captions, posts, or articles), and metadata. Drawn from fact-checking archives, the dataset reflects authentic events and may include sensitive material. For each case, participants must analyze the content and submit an English verification report summarizing findings, content categories, forensic analysis, and key evidence including source, location, time, and involved entities.
3.1.2. Sub-task
We use the COSMOS dataset (Aneja et al., 2021) for our experiments. COSMOS is a recently compiled benchmark for Cheapfakes detection. It consists of images and captions scraped from news articles and other websites, designed to capture the manipulation or misuse of multimedia content by altering the context of an image or video with misleading captions or claims. COSMOS contains 161,752 images for training, 41,006 images for validation, and 1,000 images for testing. Each image is associated with multiple captions, some of which are in-context and some are out-of-context. The test set is manually annotated with binary labels indicating whether a given image-caption pair is in-context or out-of-context.
3.2. Sub-task Metrics
To evaluate the effectiveness of the proposed method, the challenge use accuracy as the main criterion. The formula for calculating accuracy is as follows:
$accuracy=\dfrac{TP+TN}{TP+TN+FP+FN}$
where $TP$ is the number of samples correctly classified as OOC (out-of-context), $TN$ is the number of samples correctly classified as NOOC (not-out-of-context), $FP$ is the number of samples misclassified as OOC and $FN$ is the number of samples misclassified as NOOC.
3.3. Results
In this section, we present the results of our experiments on the cheapfake detection task using accuracy as the metric. We compare our proposed model with several methods on the public test of the COSMOS dataset.
3.3.1. Main-task
Table 1. Leaderboard (Total Accuracy Scores) for the Grand Challenge
Team 1 Team 2 Team 3 Ours
Score 844.22 487.19 295.86 644.65
As shown in Table 1, our system (Ours) achieved a total accuracy of 644.65, securing a clear lead over other participating teams. The margin of more than 150 points compared to the next-highest competitor demonstrates the robustness and adaptability of our verification pipeline across diverse evaluation cases. These results highlight the effectiveness of our approach in maintaining high performance throughout the challenge and its strong competitiveness in the task setting.
3.3.2. Sub-task
Table 2. Sub-task’s Performance Comparisons
Method Accuracy(%) Method Accuracy(%)
Baseline (Aneja et al., 2021) 82.0
Phan et al. (Phan et al., 2025) 83.5 Pham et al. (Pham et al., 2024) 88.9
Vo et al. (Vo-Hoang et al., 2024) 85.9 Wu et al. (Wu et al., 2023) 89.4
La et al. (La et al., 2022b) 86.7 HierOOC 95.3
Tran et al. (Tran et al., 2022) 89.1 SearchOOC 96.0
The baseline method proposed by Aneja et al. (Aneja et al., 2021) achieved an accuracy of 82.0%. Tran et al. (Tran et al., 2022) and Wu et al. (Wu et al., 2023) improved upon this result, achieving accuracies of 89.1% and 89.4%, respectively. The multimodal approach by La et al. (La et al., 2022b) reached an accuracy of 86.7%, whereas their boosting-based method (La et al., 2022a) resulted in a lower accuracy of 76.0%. The generative method proposed by Le et al. (Le et al., 2024) obtained an accuracy of 79.4%, indicating that generative approaches remain less effective for this task.
Our proposed methods, HierOOC and SearchOOC, achieved the highest accuracy scores of 95.3% and 96.0%, respectively, surpassing all existing approaches. These results demonstrate the effectiveness and robustness of our methods in the evaluated setting. In addition to accuracy, further analysis of computational efficiency and model interpretability can provide a more comprehensive evaluation of their practical advantages.
4. Conclusion
In this paper, we introduced a unified and comprehensive system for multimedia authenticity and context assessment, developed to address the pressing challenges of misinformation and disinformation on social media platforms. Our system tackles the complex task of verifying content across diverse modalities including images, videos, and text, in multilingual environments. Key contributions include a multi-stage pipeline integrating visual forensics, textual analysis, and robust multimodal reasoning. We specifically highlighted our novel hybrid approaches, SearchOOC and HierOOC, for effectively detecting out-of-context media by combining semantic similarity, temporal alignment, and geolocation cues.
Extensive evaluations on the ACM Multimedia 2025 Grand Challenge benchmark (Dang-Nguyen et al., 2025), particularly for the challenging out-of-context sub-task, demonstrate the strong effectiveness of our proposed methods. By generating both detailed expert-oriented verification reports and accessible public summaries, our system provides practical, interpretable tools vital for journalists, fact-checkers, and researchers. This work significantly advances the state of the art in multimedia verification, paving the way for more reliable information ecosystems and enhancing public trust in the digital age. Future work will focus on continuous improvement of cross-modal alignment, enhancing robustness against novel manipulation techniques, and expanding evaluation across a wider array of real-world datasets and languages.
Acknowledgements.
The authors gratefully acknowledge the financial support provided by FPT Software for covering the research registration fee associated with this paper. All intellectual property rights arising from the work shall belong solely and exclusively to the National Institute of Information and Communications Technology (NICT), and may be used by NICT for any purpose, including but not limited to commercial applications, without any claim, restriction, or obligation from or to non-NICT contributors.